# Low-Rank Graph Contrastive Learning for Node Classification

### v1

@@ Abstract
Graph Neural Networks (GNNs) have been widely used to learn node representations and with outstanding performance on various tasks such as node classification. However, noise, which inevitably exists in real-world graph data, would considerably degrade the performance of GNNs revealed by recent studies. In this work, we propose a novel and robust GNN encoder, Low-Rank Graph Contrastive Learning (LR-GCL). Our method performs transductive node classification in two steps. First, a low-rank GCL encoder named LR-GCL is trained by prototypical contrastive learning with low-rank regularization. Next, using the features produced by LR-GCL, a linear transductive classification algorithm is used to classify the unlabeled nodes in the graph. Our LR-GCL is inspired by the low frequency property of the graph data and its labels, and it is also theoretically motivated by our sharp generalization bound for transductive learning. To the best of our knowledge, our theoretical result is among the first to theoretically demonstrate the advantage of low-rank learning in graph contrastive learning supported by strong empirical performance. Extensive experiments on public benchmarks demonstrate the superior performance of LR-GCL and the robustness of the learned node representations. The code of LR-GCL is available at https://anonymous.4open.science/r/Low-Rank_Graph_Contrastive_Learning-64A6/.
@@ 1 Introduction
Graph Neural Networks (GNNs) have become popular tools for node representation learning in recent years [Kipf and Welling, 2017, Bruna et al., 2014, Hamilton et al., 2017, Xu et al., 2019]. Most prevailing GNNs  [Kipf and Welling, 2017, Zhu and Koniusz, 2020] leverage the graph structure and obtain the representation of nodes in a graph by utilizing the features of their connected nodes. Benefiting from such propagation mechanism, node representations obtained by GNN encoders have demonstrated superior performance on various downstream tasks such as semi-supervised node classification and node clustering.
Although GNNs have achieved great success in node representation learning, many existing GNN approaches do not consider the noise in the input graph. In fact, noise inherently exists in the graph data for many real-world applications. Such noise may be present in node attributes or node labels, which forms two types of noise, attribute noise and label noise. Recent works, such as [Patrini et al., 2017], have evidenced that noisy inputs hurt the generalization capability of neural networks. Moreover, noise in a subset of the graph data can easily propagate through the graph topology to corrupt the remaining nodes in the graph data. Nodes that are corrupted by noise or falsely labeled would adversely affect the representation learning of themselves and their neighbors.
While manual data cleaning and labeling could be remedies to the consequence of noise, they are expensive processes and difficult to scale, thus not able to handle almost infinite amount of noisy data online. Therefore, it is crucial to design a robust GNN encoder that could make use of noisy training data while circumventing the adverse effect of noise. In this paper, we propose a novel GCL encoder termed Low-Rank Graph Contrastive Learning (LR-GCL) to improve the robustness and the generalization capabilities of node representations for GNNs.
@@ 1.1 Contributions
First, we present a novel and provable GCL encoder termed Low-Rank Graph Contrastive Learning (LR-GCL). Our algorithm is inspired by the low frequency property illustrated in Figure 1. That is, the low-rank projection of the ground truth clean labels possesses the majority of the information of the clean labels, and projection of the label noise is mostly uniform over all the eigenvectors of a kernel matrix used in classification. Inspired by this observation, LR-GCL adds the truncated nuclear norm as a low-rank regularization term in the loss function of the regular prototypical graph contrastive learning. As a result, the features produced by LR-GCL tend to be low-rank, and such low-rank features are the input to the linear transductive classification algorithm. We provide a novel generalization bound for the test loss on the unlabeled data, and our bound is among the first few works which exhibit the advantage of learning with low-rank features for transductive classification with the presence of noise.
Second, we provide strong theoretical guarantee on the generalization capability of the linear transductive algorithm with the low-rank features produced by LR-GCL as the input. Extensive experimental results on popular graph datasets evidence the advantage of LR-GCL over competing methods for node classification on noisy graph data.
Figure 1: Eigen-projection (first row) and signal concentration ratio (second row) on Cora, Citeseer, and Pubmed. To compute the eigen-projection, we first calculate the eigenvectors $\mathbf{U}$ of the feature gram matrix $\mathbf{K}=\mathbf{H}\mathbf{H}^{\top}$, then the eigen-projection value is computed by $\mathbf{p}=\frac{1}{C}\sum_{c=1}^{C}{\left\|{\mathbf{U}}^{\top}\tilde{}\mathbf{Y}^{(c)}\right\|}_{2}^{2}/{\left\|\tilde{}\mathbf{Y}^{(c)}\right\|}_{2}^{2}\in\mathbb{R}^{N}$, where $C$ is the number of classes, and ~$\tilde{}\mathbf{Y}\in\{0,1\}^{N\times C}$ is the one-hot clean labels of all the nodes, ~$\tilde{}\mathbf{Y}^{(c)}$ is the $c$-th column of ~$\tilde{}\mathbf{Y}$. With the presence of label noise $\mathbf{N}\in\mathbb{R}^{N\times C}$, the observed label matrix is $\mathbf{Y}=\tilde{}\mathbf{Y}+\mathbf{N}$. The eigen-projection $\mathbf{p}_{r}$ for $r\in[N]$ reflects the amount of the signal projected onto the $r$-th eigenvector of $\mathbf{K}$, and the signal concentration ratio of a rank $r$ reflects the proportion of signal projected onto the top $r$ eigenvectors of $\mathbf{K}$. The signal concentration ratio for rank $r$ is computed by ${\left\|\mathbf{p}^{(1:r)}\right\|}_{2}$, where $\mathbf{p}^{(1:r)}$ contains the first $r$ elements of $\mathbf{p}$. It is observed from the red curves in the first row that the projection of the ground truth clean labels mostly concentrates on the top eigenvectors of $\mathbf{K}$. On the other hand, the projection of label noise, computed by $\frac{1}{C}\sum_{c=1}^{C}{\left\|{\mathbf{U}}^{\top}\mathbf{N}^{(c)}\right\|}_{2}^{2}/{\left\|\mathbf{Y}^{(c)}\right\|}_{2}^{2}\in\mathbb{R}^{N}$, is relatively uniform over all the eigenvectors, as illustrated by the blue curves in the first row. For example, by the rank $r=0.2\min\left\{N,d\right\}$, the signal concentration ratio of ~$\tilde{}\mathbf{Y}$ for Cora, Citeseer, and Pubmed are $0.844$, $0.809$, and $0.784$ respectively. We refer to such property as the low frequency property, which suggests that we can learn a low-rank portion of the observed label $\mathbf{Y}$ which covers most information in the ground truth clean label while only learning a small portion of the label noise. Figure 3 in the supplementary further illustrates the low frequency property on more datasets.
@@ 2 Related Works
2.1 Graph Neural Networks
Graph neural networks (GNNs) have recently become popular tools for node representation learning. Given the difference in the convolution domain, current GNNs fall into two classes. The first class features spectral convolution [Bruna et al., 2014, Kipf and Welling, 2017], and the second class [Hamilton et al., 2017, Veličković et al., 2017, Xu et al., 2019] generates node representations by sampling and propagating features from their neighborhood. To learn node representation without node labels, contrastive learning has recently been applied to the training of GNNs [Suresh et al., 2021, Thakoor et al., 2021, Wang et al., 2022, Lee et al., 2022, Feng et al., 2022a, Zhang et al., 2023, Lin et al., 2023]. Most proposed graph contrastive learning methods [Veličković et al., 2019, Sun et al., 2019, Hu et al., 2019, Jiao et al., 2020, Peng et al., 2020, You et al., 2021, Jin et al., 2021, Mo et al., 2022] create multiple views of the unlabeled input graph and maximize agreement between the node representations of these views. For example, SFA [Zhang et al., 2023] manipulates the spectrum of the node embeddings to construct augmented views in graph contrastive learning. In addition to constructing node-wise augmented views, recent works [Xu et al., 2021, Guo et al., 2022, Li et al., 2021] propose to perform contrastive learning between node representations and semantic prototype representations [Snell et al., 2017, Arik and Pfister, 2020, Allen et al., 2019, Xu et al., 2020] to encode the global semantics information.
However, as pointed out by [Dai et al., 2021], the performance of GNNs can be easily degraded by noisy training data [NT et al., 2019]. Moreover, the adverse effects of noise in a subset of nodes can be exaggerated by being propagated to the remaining nodes through the network structure, exacerbating the negative impact of noise. Unlike previous GCL methods, we propose using contrastive learning to train GNN encoders that are robust to noise existing in the labels and attributes of nodes.
2.2 Existing Methods Handing Noisy Data
Previous works [Zhang et al., 2021] have shown that deep neural networks usually generalize badly when trained on input with noise. Existing literature on robust learning mostly fall into two categories. The first category [Patrini et al., 2017, Goldberger and Ben-Reuven, 2016] mitigates the effects of noisy inputs by correcting the computation of loss function, known as loss corruption. The second category aims to select clean samples from noisy inputs for the training [Malach and Shalev-Shwartz, 2017, Jiang et al., 2018, Yu et al., 2019, Li et al., 2020, Han et al., 2018], known as sample selection. To improve the performance of GNNs on graph data with noise, NRGNN[Dai et al., 2021] first introduces a graph edge predictor to predict missing links for connecting unlabeled nodes with labeled nodes. RTGNN [Qian et al., 2022] trains a robust GNN classifier with scarce and noisy node labels. It first classifies labeled nodes into clean and noisy ones and adopts reinforcement supervision to correct noisy labels. To improve the robustness of the node classifier on the dynamic graph, GraphSS [Zhuang and Al Hasan, 2022] proposes to generalize noisy supervision as a kind of self-supervised learning method, which regards the noisy labels, including both manual-annotated labels and auto-generated labels, as one kind of self-information for each node. Different from previous works, we aim to improve the robustness of GNN encoders for node classification by applying low-rank regularization during the training of the transductive classifier.
@@ 3.1 Notations
An attributed graph consisting of $N$ nodes is formally represented by $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denote the set of nodes and edges respectively. $\mathbf{X}\in\mathbb{R}^{N\times D}$ are the node attributes, and the attributes of each node is in $\mathbb{R}^{D}$. Let $\mathbf{A}\in\{0,1\}^{N\times N}$ be the adjacency matrix of graph $\mathcal{G}$, with $\mathbf{A}_{ij}=1$ if and only if $(v_{i},v_{j})\in\mathcal{E}$. $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ denotes the adjacency matrix for a graph with self-loops added. $\tilde{\mathbf{D}}$ denotes the diagonal degree matrix of $\tilde{\mathbf{A}}$. $[n]$ denotes all the natural numbers between $1$ and $N$ inclusively. ${\cal L}$ is a subset of $[N]$ of size $m$, and ${\cal U}$ is a subset of $[N]\setminus{\cal L}$ and $\left|{\cal U}\right|=u$. Let ${\cal V}_{{\cal L}}$ and ${\cal V}_{{\cal U}}$ denote the set of labeled nodes and unlabeled test nodes respectively, and $\left|{\cal V}_{{\cal L}}\right|=m$, $\left|{\cal V}_{{\cal U}}\right|=u$. Note that $m+u\leq N$, and it is not necessary that $m+u=N$ because there are usually validation nodes other than the labeled nodes and unlabeled test nodes. Let $\mathbf{u}\in\mathbb{R}^{N}$ be a vector, we use $\left[\mathbf{u}\right]_{{\cal A}}$ to denote a vector formed by elements of $\mathbf{u}$ with indices in ${\cal L}$ for ${\cal A}\subseteq[N]$ If $\mathbf{u}$ is a matrix, then $\left[\mathbf{u}\right]_{{\cal A}}$ denotes a submatrix formed by rows of $\mathbf{u}$ with row indices in ${\cal A}$. ${\left\|\cdot\right\|}_{\textup{F}}$ denotes the Frobenius norm of a matrix, and ${\left\|\cdot\right\|}_{p}$ denotes the $p$-norm of a vector.
3.2 Graph Convolution Network (GCN)
@@
To learn the node representation from the attributes $\mathbf{X}$ and the graph structure $\mathbf{A}$, one simple yet effective neural network model is Graph Convolution Network (GCN). GCN is originally proposed for semi-supervised node classification, which consists of two graph convolution layers. In our work, we use GCN as the backbone of the proposed LR-GCL, which is the GCL encoder, to obtain node representation $\mathbf{H}\in\mathbb{R}^{N\times d}$, where the $i$-th row of $\mathbf{H}$ is the node representation of $v_{i}$. In this manner, the output of LR-GCL is $\mathbf{H}=g(\mathbf{X},\mathbf{A})=\sigma(\widehat{\mathbf{A}}\sigma(\widehat{\mathbf{A}}\mathbf{X}{\tilde{}\mathbf{W}}^{(0)}){\tilde{}\mathbf{W}}^{(1)})$, where $\widehat{\mathbf{A}}={\tilde{\mathbf{D}}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}$, ~${\tilde{}\mathbf{W}}^{(0)}$ and ~${\tilde{}\mathbf{W}}^{(1)}$ are the weight matrices, and $\sigma$ is the activation function ReLU. The robust and low-rank node representations produced by the LR-GCL are used to perform transductive node classification by a linear classifier. LR-GCL and the linear transductive node classification algorithm are detailed in Section 4.

### v2

# Graph Contrastive Learning with Low-Rank Regularization and Low-Rank Attention for Noisy Node Classification

@@ Abstract
Graph Neural Networks (GNNs) have achieved remarkable success in learning node representations and have shown strong performance in tasks such as node classification. However, recent findings indicate that the presence of noise in real-world graph data can substantially impair the effectiveness of GNNs. To address this challenge, we introduce a robust and innovative node representation learning method named Graph Contrastive Learning with Low-Rank Regularization, or GCL-LRR, which follows a two-stage transductive learning framework for node classification. In the first stage, the GCL-LRR encoder is optimized through prototypical contrastive learning while incorporating a low-rank regularization objective. In the second stage, the representations generated by GCL-LRR are employed by a linear transductive classifier to predict the labels of unlabeled nodes within the graph. Our GCL-LRR is inspired by the Low Frequency Property (LFP) of the graph data and its labels, and it is also theoretically motivated by our sharp generalization bound for transductive learning. To the best of our knowledge, our theoretical result is among the first to theoretically demonstrate the advantage of low-rank regularization in transductive learning, which is also supported by strong empirical results. To further enhance the performance of GCL-LRR, we present an improved model named GCL-LR-Attention, which incorporates a novel LR-Attention layer into GCL-LRR. GCL-LR-Attention reduces the kernel complexity of GCL-LRR and contributes to a tighter generalization bound, leading to improved performance. Extensive evaluations on standard benchmark datasets evidence the effectiveness and robustness of both GCL-LRR and GCL-LR-Attention in learning meaningful node representations. The code is available at https://github.com/Statistical-Deep-Learning/GCL-LR-Attention.
keywords:
Graph Neural Networks, Contrastive Learning, Low-Rank Regularization, Noisy Node Classification
@@ 1 Introduction
Graph Neural Networks (GNNs) are widely recognized as effective tools for node representation learning, with numerous models [1, 2, 3, 4] demonstrating success across a range of tasks. Among them, methods such as [1, 5] utilize the graph topology by aggregating information from neighboring nodes. This form of feature propagation enables the models to generate powerful embeddings, leading to strong outcomes in semi-supervised node classification and clustering. However, the majority of existing GNN methods do not adequately address the presence of noise in the graph data [6, 7]. Such noise may arise in the attributes or labels of nodes, introducing attribute noise and label noise, respectively. Prior studies [8] have demonstrated that noise in the input data can significantly impair the generalization ability of neural networks. In the case of graphs, this issue is further amplified since noise associated with a few nodes can spread through the graph structure and affect other nodes [9, 10, 11]. As a result, corrupted nodes not only degrade their own representations but also influence those of their neighbors. Although it is possible to manually clean or relabel data, such approaches are costly and do not scale well. This challenge underscores the need for GNN models that are capable of learning effectively even in the presence of noisy inputs.
To this end, we introduce Graph Contrastive Learning with Low-Rank Regularization, abbreviated as GCL-LRR, which is a new encoder aimed at improving both robustness and generalization in node representation learning. Traditional strategies for robust learning generally fall into two categories. The first modifies the loss function to accommodate corrupted data [8, 12], while the second removes samples suspected to be noisy [13, 14, 15, 16, 17]. Although some of these ideas have been adapted for graph settings [9, 18, 19], they often depend on heuristics and lack theoretical backing in semi-supervised scenarios. The design of our GCL-LRR model is inspired by the low-frequency nature of graph signals, and it is also theoretically supported by a new generalization bound developed for transductive learning. To the best of our knowledge, this paper is among the first to provide a principled justification for the advantage of low-rank representation learning in graph contrastive settings. Experimental evaluations conducted on widely used benchmarks demonstrate that GCL-LRR consistently achieves strong robustness and high performance. Although GNNs are known to function as low-pass filters, they do not explicitly target low-frequency signals. Consequently, their ability to exploit the Low-Frequency Property (LFP) in noisy labels remains limited. As visualized in Figure 1, the LFP reveals that clean label information tends to concentrate within the low-rank structure of the observed label space. Unlike conventional GNNs, the GCL-LRR encoder explicitly captures this structure by learning representations constrained to be low-rank. Prior work [20] has illustrated the benefit of such low-rank learning in mitigating attribute noise by introducing learnable filtering mechanisms. Moreover, recent studies involving graph attention and transformer-based architectures emphasize the need to balance both low- and high-frequency components for improved node representations [21, 22]. In comparison, GCL-LRR offers an improved balance by promoting low-frequency information through minimization of the Truncated Nuclear Norm (TNN), which aligns with the LFP. Our GCL-LRR method is also grounded in theory through our novel generalization bound in the transductive setting. To further boost the performance of GCL-LRR, we introduce an enhanced model termed GCL-LR-Attention, which integrates a novel LR-Attention layer into GCL-LRR. GCL-LR-Attention lowers the kernel complexity of GCL-LRR by the LR-Attention layer and yields a tighter generalization error bound than GCL-LRR, leading to even better performance than GCL-LRR. Performance comparisons in Table 3 of Section 5.2 demonstrate that GCL-LRR and GCL-LR-Attention surpass state-of-the-art methods based on attention and transformers, such as GFSA [21] and HONGAT [22], when evaluated under label and attribute noise. Furthermore, the trade-off between low- and high-frequency learning is evaluated through kernel complexity, a metric detailed in Section 4.2. As shown in Table 4 of Section 5.3, the GCL-LRR model achieves lower kernel complexity, leading to a lower upper bound on the test loss than competing graph contrastive and attention-based methods, confirming the effectiveness of the learned low-rank representations. Benefiting from the LR-Attention layer, GCL-LR-Attention achieves an even lower generalization bound and kernel complexity, leading to even better node classification performance under various types of noise.
@@ 1.1 Contributions
First, we introduce Graph Contrastive Learning with Low-Rank Regularization (GCL-LRR), a novel and theoretically grounded GCL encoder. The design of GCL-LRR is motivated by the LFP illustrated in Figure 1, which shows that a low-rank projection of the clean label matrix captures most of its informative content. In contrast, label noise tends to spread uniformly across all eigenvectors of the classification kernel matrix. Drawing from this insight, GCL-LRR incorporates the TNN as a low-rank regularization component into the loss function of standard prototypical graph contrastive learning. This regularization encourages the model to produce inherently low-rank representations, which are subsequently used in a linear transductive classifier.
Second, we establish a strong theoretical foundation that guarantees the generalization performance of the linear transductive algorithm when applied to the low-rank features generated by GCL-LRR. In particular, we derive a novel generalization bound on the test loss for unlabeled nodes. To the best of our knowledge, this theoretical result is among the first demonstrating the benefit of learning low-rank representations for robust transductive classification under label noise. Motivated by this theoretical result and in pursuit of further performance gains, we introduce an enhanced variant of our model that incorporates a novel LR-Attention layer, termed the GCL-LR-Attention. GCL-LR-Attention achieves a further reduction in kernel complexity compared to GCL-LRR, which leads to a tighter bound on the generalization error in the transductive setting and improved empirical performance. As demonstrated in Table 4 of Section 5.3, the GCL-LRR model has a lower generalization upper bound than existing methods. The GCL-LR-Attention model further reduces the generalization upper bound of the GCL-LRR model. Comprehensive experiments conducted on widely used graph benchmarks demonstrate the superiority of both GCL-LRR and GCL-LR-Attention over existing methods in node classification tasks involving noisy graph data.
@@ 2 Related Works
2.1 Graph Neural Network and its Training on Noisy Data
Graph neural networks (GNNs) have established themselves as effective approaches for learning node representations. Existing GNNs are generally grouped into spectral methods [2, 1] and spatial methods [3, 23, 4], both of which learn by aggregating information from the local neighborhoods of nodes. In recent years, contrastive learning has become increasingly prominent for unsupervised learning on graphs [24, 25, 26, 27, 28, 29, 30]. A majority of graph contrastive learning (GCL) methods [31, 32, 33, 34, 35, 36, 37, 38] generate multiple augmented views of the original graph and encourage the alignment between the embeddings corresponding to each view. In addition to augmenting at the node level, some recent efforts [39, 40, 41] have improved GCL by introducing semantic prototypes [42, 43, 44, 45] and using these prototypes to contrast node features against global representations.
A major challenge in training deep neural networks, including GNNs, lies in their vulnerability to noisy inputs, which has been highlighted in earlier research [46]. Work on robust learning generally follows two principal strategies. One approach adjusts the objective function during training to account for corrupted data, a technique commonly known as loss correction [8, 12]. Another approach focuses on identifying clean samples from noisy datasets and using them for training, a process referred to as sample selection [13, 14, 15, 16, 17]. Within the domain of graph-based learning, recent contributions aim to improve robustness by incorporating techniques such as structural prediction, denoising of labels, and self-supervised objectives [9, 18, 19, 47, 48]. Departing from these earlier directions, our approach focuses on enhancing the robustness of GNN encoders for node classification by embedding low-rank regularization directly into the training process of the GCL encoder.
2.2 Learning Low-Frequency Signal in Graphs with GNNs and Graph Attention
Traditional graph neural networks, such as GCN [1], operate by aggregating information from neighboring nodes and therefore naturally function as low-pass filters. Earlier investigations [49, 50, 51, 52] underscore the significance of capturing low-frequency signals that are inherent in both graph structures and node features. Nevertheless, an exclusive reliance on low-frequency components can introduce the problem of over-smoothing [53, 54, 55, 56], which becomes particularly detrimental in scenarios where inter-class links are abundant. To address this issue, recent studies [53, 57, 58] have introduced adaptive techniques that aim to balance low- and high-frequency signals, achieving better results in node classification tasks [59]. Alongside these developments, models that prioritize learning from low-rank components of graph features and structures have been shown to be more robust in the presence of noise [60, 61]. Graph attention mechanisms such as GAT [23] are also capable of emphasizing low-frequency patterns [22, 62]. For instance, HONGAT [22] combats the over-smoothing effect by enhancing high-order dependencies and introducing sparsity into the attention weights. Additionally, combining spectral filtering with attention mechanisms has emerged as a promising direction to enable adaptive frequency-aware learning in node representation models [63, 64, 65].
@@ 3.1 Notations
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$ denote an attributed graph with $N$ nodes. The node set is given by $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ and the edge set satisfies $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. The matrix $\mathbf{X}\in\mathbb{R}^{N\times D}$ contains the attribute information of all nodes, where $D$ corresponds to the dimensionality of each node’s attributes. The adjacency matrix associated with $\mathcal{G}$ is denoted by $\mathbf{A}\in\{0,1\}^{N\times N}$ and satisfies $\mathbf{A}_{ij}=1$ whenever there is an edge $(v_{i},v_{j})$ in $\mathcal{E}$. When self-loops are added to the graph, the resulting adjacency matrix becomes $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$, and the corresponding degree matrix is defined as $\tilde{\mathbf{D}}$, which is diagonal. The notation $[N]$ refers to the set of integers from $1$ to $N$ inclusive. A subset ${\cal L}\subseteq[N]$ contains $m$ labeled nodes, and its complement ${\cal U}=[N]\setminus{\cal L}$ has cardinality $u$. The sets ${\cal V}_{{\cal L}}$ and ${\cal V}_{{\cal U}}$ represent the collections of labeled and unlabeled nodes respectively, with $\left|{\cal V}_{{\cal L}}\right|=m$ and $\left|{\cal V}_{{\cal U}}\right|=u$. For any vector $\mathbf{u}\in\mathbb{R}^{N}$, the notation $\left[\mathbf{u}\right]_{{\cal A}}$ refers to the subvector composed of entries indexed by ${\cal A}\subseteq[N]$. In the case where $\mathbf{u}$ is a matrix, $\left[\mathbf{u}\right]_{{\cal A}}$ denotes the submatrix consisting of the rows indexed by ${\cal A}$. The Frobenius norm of a matrix is denoted by ${\left\|\cdot\right\|}_{\textup{F}}$, and the $p$-norm of a vector is expressed as ${\left\|\cdot\right\|}_{p}$.
3.2 Graph Convolution Network
@@
A straightforward yet effective approach for learning node representations from the attribute matrix $\mathbf{X}$ and the graph structure $\mathbf{A}$ is the Graph Convolutional Network (GCN). Originally introduced for semi-supervised node classification, GCN is composed of two layers of graph convolution operations. In our framework, we adopt GCN as the base architecture within the proposed GCL-LRR, which serves as the GCL encoder, to generate node embeddings denoted by ^$\widehat{}\mathbf{H}\in\mathbb{R}^{N\times d}$. Each row ^$\widehat{}\mathbf{H}_{i}$ corresponds to the representation of node $v_{i}$. The computation in GCL-LRR follows the formulation ^$\widehat{}\mathbf{H}=g(\mathbf{X},\mathbf{A})=\sigma(\widehat{\mathbf{A}}\sigma(\widehat{\mathbf{A}}\mathbf{X}{\tilde{}\mathbf{W}}^{(0)}){\tilde{}\mathbf{W}}^{(1)})$, where the matrix $\widehat{\mathbf{A}}$ is defined as ${\tilde{\mathbf{D}}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}$, and the parameters ~${\tilde{}\mathbf{W}}^{(0)}$ and ~${\tilde{}\mathbf{W}}^{(1)}$ are learnable weight matrices. The non-linear transformation $\sigma$ represents the ReLU activation function. The resulting node representations produced by GCL-LRR are both robust and low-rank, and they are subsequently used as input to a linear classifier for transductive node classification. Further details regarding the GCL-LRR encoder and the linear transductive classification procedure are provided in Section 4.